The neighborhood children have been cutting diagonally across her lawn instead of walking around the yard. If the diagonally distance across the lawn is 30 feet and the longer part of the sidewalk is twice the shorter lenght, how many feet, are the children saving by cutting across the lawn?
Answer by mananth(16949)

The neighborhood children have been cutting diagonally across her lawn instead of walking around the yard. If the diagonally distance across the lawn is 30 feet and the longer part of the sidewalk is twice the shorter lenght, how many feet, are the children saving by cutting across the lawn?
...
diagonal = 30 feet
..
on side = x
other side 2x
...
right triangle
x^2+(2x)^2=30^2
x^2+4x^2= 900
5x^2=900
/5
x^2= 180
..
x= +/- sqrt(180)
x= +13.4 feet one side
other side = 2* 13.4 = 26 .8 feet
...
sum of sides = 13.4 +26.8 = 40.2 feet
hypotenuse = 30
difference = 40.2-30=10.2 feet is waht they save.
...
m.ananth@hotmail.ca